A triangle drawn on a map has sides of lengths 8 cm, 12 cm, and 13 cm. The shortest of the corresponding real-life distances is 120 km. Find the longest of the real-life distances?
Please help me with this.

Answer by edjones(8007)

x/13=120/8
x=1560/8
=195 km
.
Ed